OpenStudy (anonymous):

how do you solve 12=-4(-6x-3)

sam (.sam.):

12=-4(-6x-3) -4(-6x-3)=12 -(-6x-3)=3 6x+3=3 6x=0 x=0

OpenStudy (anonymous):

You first solve the parenthesis towards the right side of the '=' This will give you your first step, i.e. 1) 12 = (-4*-6x) - (3*-4) [remember that two negatives multiplied give a positive] 2) 12 = 24x +12 [you now move the '12' on the right side to the left side of the equation] 3] 12 -12 = 24x 4) 0 = 24x x= 0 Answer